# SPSysML: A meta-model for quantitative evaluation of Simulation-Physical Systems

### v3

@@ Abstract
Robotic systems are cyber-physical systems (CPS) equipped with multiple sensors and effectors, making them aware of dynamic environments, e.g. surrounding humans. A multi-component structure and potential environmental harm make robotic systems development challenging. The Digital Twin (DT) concept has recently been applied to model and test CPS. However, it is unclear how to employ DT in robotic system development, e.g. in-development testing and how to design systems featuring various simulated and physical parts.
Goal: We aim to improve the integration between simulated and physical parts of CPS in various system setups, resulting in controlling the simulated and real hardware with a shared controller. The goal for tightening the integration is to allow parallel and safer system development and faster and safer machine learning using simulation/DTs. We propose integrity evaluation with quantitative factors and requirement-based system structure design, allowing DT to perceive exogenous actions in the simulated world.
Method: We propose a Domain-Specific Language based on Systems Modeling Language (SysML) called SPSysML. SPSysML defines the taxonomy of a CPS consisting of at least a physical or simulated part. The simulated parts can be DTs. We propose a development procedure called SPSysDP that manages the considered systems development by evaluating the system designs with the proposed quantitative factors.
Result: SPSysDP is used to develop a robotic system for the INCARE project. In subsequent iterations of SPSysDP, the simulation-physical integrity of the system is improved, and a more significant fraction of the system components is shared between various system setups. The designed system was deployed on the physical robot and two simulators. System setups are based either on Robot Operating System (ROS) or ROS2. Therefore, we argue that SPSysDP and SPSysML are neither specific for a control system framework nor a robot simulator.
Conclusion: SPSysML with SPSysDP enables the design of SPSys (including DT and CPS), multi-setup system development featuring improved integrity between simulation and physical parts in its setups.
@@ 1 Introduction
Simulation is widely used in state-of-the-art development procedures for cyber-physical systems (CPS). Recent papers refer to Model-in-the-loop (MIL) [1], Software-in-the-loop (SIL) [2], Hardware-in-the-loop (HIL) [3] and Rapid Control Prototyping (RCP) [4] techniques. They are used selectively or are composed in sequence, e.g. verification and validation steps in a systems development approach called V-model [5]. Each of the techniques requires a simulation of the system parts. Some systems utilise the Digital Twin (DT) [6, 7] concept. They employ an accurate simulation of a system part for, e.g.:
  • Swapping a malfunctioned system part with the simulated one DT that mirrors the part’s functionality in the simulated world,
  • Energy consumption analysis,
  • System failure analysis and prediction,
  • Technology integration,
  • Real-time monitoring.
@@
Some parts (software and hardware) of such a system interact with the real world, while other parts interact with the simulated world. We call the set of the parts used in the real world the physical embodiment, and we call the set of the parts used in the simulated world the simulated embodiment. The Digital Twin is a part of the simulated embodiment mirroring a part of the physical embodiment. DT concept is used in numerous domains and applications [8]. They are systematically analysed by the authors of [9].
@@
Robots comprise multiple devices and their controllers. In particular, autonomous mobile robots require a complex navigation system that features multiple closed control loops, e.g. drive controllers, trajectory controllers, and Simultaneous Localisation and Mapping (SLAM). Furthermore, the navigation system and the rest of the robot’s functionality are used to execute the user request (like object transportation). In some cases, only a fraction of such a complex system must have DTs. In other cases, if the robot system has only the simulated embodiment, it is a demonstrator of the future product.
@@
To clear up the taxonomy of systems featuring simulated and physical parts, we introduce the Simulation-Physical System (SPSys) concept (Fig. 1). This kind of system consists of at least one physical/simulated embodiment and a shared controller. If it has both embodiments, the simulated part can be a DT of a physical part. If it has only the physical embodiment, it is a CPS; if it has only the simulated embodiment, it is a simulator.
Figure 1: The idea of Simulation-Physical System
@@
In the development process, the system parts evolve from simulated mockups to the physical parts and the software deployed on the hardware. The system development procedure must guide the project team through the process. The system design created as a milestone in the procedure must specify all required parts in all development stages and compose them in the testing and deployment setups. Developing reliable systems, especially complex ones like robot systems, requires comprehensive unit and integration testing. Some CPS parts can be tested with simulated hardware only; therefore, additional parts that do not comprise the operational system are required (e.g. human simulator in a social robot case). Comprehensive testing is complex in test case specification and time-consuming in test implementation and execution. From this perspective, software reusability is a key to fast development of complex, reliable systems.
@@
  • Specify profiled-based requirements of Simulation-Physical Systems including multiple Digital Twins for procedural system structuring based on the requirement types,
@@
  • Re-use of software between DT and its Physical Twin (PT) and between different system setups, improving system’s reliability and resulting in more accessible and faster testing,
  • The integrity boost between simulation and physical embodiments of SPSys. This enables a comprehensive and parallel simulation-based testing of the system parts,
@@
  • Enable Digital Twins to observe exogenous actions in simulation. The actions can be executed by any agent outside the physical system, e.g. a human moving objects in a service robot environment. Such an action must be implemented in the simulated world because the digital twin must operate in the exact environment as its physical counterpart.
@@
To reach these goals, we propose a Domain-Specific Language (DSL) named Simulation-Physical System Modeling Language (SPSysML). SPSysML models the requirements and the SPSys, in particular, SPSys consisting of simulated parts (used either in the system development or in the operational setup as DT/simulator). Based on SPSysML, we propose a requirement-based SPSys structuring method. To measure the inter-embodiment integrity of the system, we propose quantitative design evaluation factors for SPSys designs. The factors are used in the proposed SPSys Development Procedure (SPSysDP) (Fig. 2) to maximise the shared controller, minimise the system parts number and improve simulation-physical integrity.
Our goal originates from the conclusion of [10]: The digital models are mainly used to examine product performance(…). However, how to optimise the use of those models to enhance the design process and design collaboration still needs to be investigated.
Figure 2: The concept of DES development procedure
SPSysML comprises a componet-based Platform Independent Model (PIM), thus, can be applied to any SPSys. The result of SPSysDP is an implemented and tested Platform Specific Model (PSM) that can be launched in the specified setups. PSM implementation utilises platform-specific tools and software libraries. Therefore, to verify our approach, we design, implement, test and deploy a specific SPSys, including the TIAGo service robot [11] for the INCARE (Integrated Solution for Innovative Elderly Care) project [^1] http://aal-incare.eu/, robot application demonstration (simulation/real) can be found accessing Demo / RAPP Store Demo. The robot features an extended voice interface and additional devices to serve the elderly, e.g. in object transportation (Fig. 2(a)) and fall assistance (Fig. 2(b)). The resulting system utilises component-based frameworks like ROS [12, 13], ROS2 [14], OROCOS [15], that are a standard open-source robot control frameworks. We use two open-source simulators Gazebo [16] (in particular Gazebo_ros_control package [17]) and O3DE [18] for DT implementation. By the SPSysDP execution for a real use-case robot system development, we present an example of DT and PT implementation in robotic systems. To make the SPSys development easier for the community, we share the SysML profiles, the SPSysML meta-model and the example INCARE system model [^2] https://github.com/RCPRG-ros-pkg/spsysml. All diagrams describing SPSysDP, SPSysML and the INCARE system are also shared [19].
@@
In this article, we describe related work concerning mixing simulation-physical setups and meta-models for CPS (Sec. 2) and state our work’s novelty and its result regarding the related works (Sec. 3). Subsequently, we define SPSys requirements model in Sec. 4.1, the meta-model in Sec. 4.2 and its application procedure in Sec. 4.3. Next, we analyse the proposed evaluation factors of SPSys design and describe features of the systems scored edge case values of the factors in Sec. 5. Finally, we verify our method in complex SPSys development for the INCARE project (Sec. 6). This work is concluded in (Sec. 7).
@@
TABLE I: Related works comparison with this work
System development method EM EIV Formal meta-model SCP SO System structure configurability DE E QE MDT Purpose
Digital Mockup [20] I - - ES - Single structure - - - RT simulation for HIL testing
C2PS [21] M - Own meta-model Formal and detailed BS Single adaptive - - N/D Digital Twin in the cloud
of work [22] M Dolev-Yao [23] Formal and detailed - Single structure - - Cyber-security, physical-digital twin synchronisation
DEVSRT [24] I - DEVS [25] General - Single structure - - - - Simulation to embedded continuous development
aRD[26] M - - General - - - - RT & NonRT parts integration
HMLF [27] I - - ES - Single structure - - - - Simulation-driven ML
RSHPN [28] M - RSHPN ES - Single structure - - - Deadlocks verification, automated code generation
SPSysDP M SPSysML /SysML Formal and detailed SPO unbounded system setups Structure optimisation for Simulation-Physical Systems
EM – Embodiment modularity, EIV – Embodiment integrity verification, SCP – System creation procedure, SO – Structure optimisation, DE – If the system execution configuration can contain parts of both embodiments, E – If uncontrolled agents of the physical environment (link humans) can be modelled in the simulation, QE – If quantitative evaluation factors for system design are proposed, MDT – If a multi Digital Twin system model is presented, I – Integrated system, ES – Example system as the method application procedure, M – System decomposable to multiple simulation/physical parts, BS – The best selection from the previously designed, SPO – Simulated-Physical design optimisation, N/D – Possible, however, not defined,
@@ 2 Related work
Our work intersects two areas of the systems engineering domain: the development of cyber-physical systems consisting of simulated parts and specifying systems following a Model-based approach. Therefore, we focus the related work analysis on development approaches for systems implementing physical parts and their simulated equivalents (sometimes called Digital Twins). The second section describes state-of-the-art Domain Specification Languages (DSL), emphasising the robotics domain. The works regard the development of various kinds of SPSys. Therefore, the purposes of their results differ. The most related works to SPSysML are compared in Table I, and their purpose is also shown.
2.1 Mixing simulation-physical setups
In [26], the authors note the complexity of robotic systems and that heterogeneous teams of researchers and developers often develop the system components simultaneously. The authors suggest a demand for system simulation enabling rapid prototyping and an iterative development process. They propose the so-called agile Robot Development (aRD) concept that facilitates the integration of the hard real-time part and the non-real-time part of the system utilising a Matlab/Simulink toolchain. The sim2real problem is known and mentioned in the recent research results, e.g. in the domain of self-driving cars [29]. However, to our knowledge, none of the existing works specifies a flexible robotic system meta-model that defines mixed simulation-physical setups of CPS under development. Agile development implements the idea of rapid prototyping via, e.g. unit testing [30, 31]. The Robot Unit Testing methodology [30] enables covering the whole robot software with unit testing, starting from the components integrating black-boxed vendor drivers. On the other hand, [31] presents a framework for test-driven development (TDD) of multi-agent systems. TDD methodology is often used in the robotics domain. However, the tests evaluate just one specified simulation-physical setup of the system. Still, during development, almost every part of the system evolves from mockup (realised in simulation) to deployable software or hardware. Continuous integration is a valuable tool for verifying a model-to-reality transformation [32]. In the CPS case, continuous integration may not be limited to software; however, it should include mechanical parts development and their integration with the software. State-of-the-art industry-implemented development operations already employ mechanical parts simulation in Gazebo for continuous integration [33]. None of the works proposes a quantitative evaluation factors for the system design assessment.
@@ 2.2 Meta-models for CPS
In the development process, a language is required to plan, conceptualise, and specify the system. Currently, the state-of-the-art approach is a model-based approach utilising a DSL. Unified Modeling Language (UML) is the most known language; however, System Modeling Language (SysML) [35] is proposed by the Object Management Group (OMG) to support the design, analysis and verification/validation of complex systems comprising software and hardware components. Different Model Driven Engineering approaches exist in the robotics domain [36]. For instance, Embodied Agent-Based cybeR-physical control systems modelling Language (EARL) [37] is SysML specialisation for robotics which allows analysis and specification of the robotic system properties. It is based on the Embodied Agent approach [38]. The approach shapes the robotic system structure and inter-component communication by defining SPSys-related artefacts, component classes and their constraints. Other DSLs support verification and testing of, e.g. industry 4.0 plants [39], or agent-based computational systems [40]). Besides, the work [41] describes the specification of safety compliance needs for critical computer-based and software-intensive systems.
2.3 Characteristics of the related works
The related models in Table I are SPSys of different types or are tools for testing/designing SPSys. Two of the works design interaction between the DTs and PTs – C2PS [21], [22], three focus on simulation-based development of CPS – Digital Mockup [20], DEVSRT [24] and aRD [26], and the last one – HMLF [27] utilises DT concept in machine learning. The authors of RSHPN [28] propose a meta-model to verify the lack of deadlocks in multi-component SPSys using Hierarchical Petri Nets. The nets are generated from a DSL-based specification, and the code associated with the robot controller structure and communication is generated.
3 The work novelty and result
@@
None of the models presented in Sec. 2 specifies SPSys in general, especially including a physical part without equivalent Digital Twin, simulated parts without equivalent PT, or hybrid SPSys including Twins and non-Twin components. The key novel features delivered by SPSysML and SPSysDP are:
@@
  • Integrity evaluation with quantitative factors,
  • SysML-based structure model and system development procedure,
  • System design optimisation considering simulation and physical embodiments of the system,
  • Multiple system execution and testing setups specification,
  • Forcing simulation-based testing prior to physical embodiment development,
@@
  • Specification of Simulation-Physical System including multiple physical and simulated parts and multiple Digital Twins,
@@
  • Reflection of the dynamic environment in simulation to enable Digital Twins to observe exogenous actions, e.g. human moving objects in the simulated world.
@@
Based on our experience in SPSys development (TIAGo robot [42], Velma robot [43], IRP6 robot [^3] Real: https://youtu.be/wJpFcy99Gh0, Simulation: https://youtu.be/BjwcbSdouHw [44, 45]) and the literature analysis, we propose the SPSys Development Procedure focused on simulation-based testing coverage and software integrity between the system embodiments maximisation. Complex system structuring is a broad topic with solutions utilising various methods and strategies, e.g. top-down, middle-out and bottom-up. In [46], we describe a binary communication-focused top-down approach for robotic systems. In this article, we propose a requirement-based top-down system structuring method customised for SPSys.
4 Simulation-Physical System

### v5

@@ Abstract
Evaluating early design concepts is crucial as it impacts quality and cost. This process is often hindered by vague and uncertain design information. This article introduces the SysML-based Simulated-Physical Systems Modeling Language (SPSysML). It is a Domain-Specification Language for evaluating component reusability in Cyber-Physical Systems incorporating Digital Twins and other simulated parts. The proposed factors assess the design quantitatively. SPSysML uses a requirement-based system structuring method to couple simulated and physical parts with requirements. SPSysML enables DTs to perceive exogenous actions in the simulated world.
SPSysML validation is survey- and application-based. First, a robotic system for an assisted living project was developed. The integrity of simulated and physical parts of the system is improved using SPSysML-based quantitative evaluation. Thus, more system components are shared between the simulated and physical setups. The system was deployed on the physical robot and two simulators based on the Robot Operating System (ROS) or ROS2. SPSysML was used by a third-party developer and was assessed by him and other practitioners in a survey.
keywords:
[…]
@@ 1 Introduction
Simulation is widely used in state-of-the-art cyber-physical systems (CPS) development procedures. Recent papers refer to Model-in-the-loop (MIL) [1], Software-in-the-loop (SIL) [2], Hardware-in-the-loop (HIL) [3] and Rapid Control Prototyping (RCP) [4] techniques. They are used selectively or are composed in sequence, e.g. verification and validation steps in a systems development approach called V-model [5]. Each of the techniques requires a simulation of the system parts. Some systems utilise the Digital Twin (DT) [6, 7] concept. They employ an accurate simulation of a system part for, e.g. Swapping malfunctioned parts, Energy consumption analysis, Technology integration, or Real-time monitoring.
@@
The system’s parts (software and hardware) interact with the real world, while others interact with the simulated world. We define the set of the parts used in the real world as the physical embodiment and the set of the parts used in the simulated world as the simulated embodiment. The DT is a part of the simulated embodiment mirroring a part of the physical embodiment. DT concept is used in numerous domains and applications [8]. They are systematically analysed by the authors of [9].
@@
Robots comprise multiple devices and their controllers. In particular, autonomous mobile robots require a complex navigation system featuring multiple closed control loops, e.g. drive controllers, trajectory controllers, and Simultaneous Localisation and Mapping (SLAM). Furthermore, in a complex mobile manipulator system [10], the navigation system is integrated with manipulation control to execute user requests (like object transportation). In some cases, only a fraction of such a complex system must have DTs. In other cases, if the robot system only has the simulated embodiment, it is a demonstrator of a future product.
@@
To clear up the taxonomy of systems featuring simulated and physical parts, we introduce the Simulated-Physical System (SPSys) concept (Fig. 1). This kind of system consists of at least one physical/simulated embodiment and a shared controller. If it has both embodiments, the simulated part can be a DT of a physical part. If it has only the physical embodiment, it is a CPS; if it has only the simulated embodiment, it is a simulator.
Figure 1: The idea of Simulated-Physical System
@@
Developing reliable systems, especially complex ones like robot systems, requires comprehensive unit and integration testing. Some CPS parts can be tested with simulated hardware only; therefore, additional parts that do not comprise the operational system are required (e.g. human simulator in a social robot case). Comprehensive testing is complex in test case specification and time-consuming in test implementation and execution. From this perspective, software reusability is a key to fast development of complex, reliable systems.
@@
  • Enhance system requirements traceability for simulated and physical system parts thanks to the introduction of profiled-based requirements of Simulated-Physical Systems, including multiple Digital Twins for procedural system structuring based on the requirement types,
@@
  • Evaluate the level of component re-use between DT and its Physical Twin (PT) and between different system setups, improving the system’s reliability and resulting in more accessible and faster testing,
@@
  • Enable DTs to observe exogenous actions in simulation. The actions can be executed by any agent outside the physical system, e.g. a human moving objects in a service robot environment. Such an action must be implemented in the simulated world to make DT operate in the exact environment as its physical counterpart.
@@
To reach these goals, we propose a Domain-Specific Language (DSL) named Simulated-Physical System Modeling Language (SPSysML). SPSysML models the requirements and the SPSys, in particular, SPSys consisting of simulated parts (used either in the system development or in the operational setup as a DT/simulator). Based on SPSysML, we propose a requirement-based SPSys structuring method for enhancing requirement readability in simulated and physical system parts. To measure the inter-embodiment integrity of the system, we propose quantitative design evaluation factors for SPSys designs. The factors are used during system development to maximise the shared controller, minimise the variety of the system parts and improve simulated-physical integrity. Our goal originates from the conclusion of [11]: The digital models are mainly used to examine product performance(…). However, how to optimise the use of those models to enhance the design process and design collaboration still needs to be investigated.
@@
In this article, we describe related work concerning mixing simulated-physical setups and meta-models for CPS (Sec. 2) and state our work’s novelty and its result regarding the related works (Sec. 3). Subsequently, we define SPSys requirements model in Sec. 4.1, the meta-model in Sec. 4.2. In Sec. 5, we analyse the proposed evaluation factors of SPSys design and describe characteristics of a system scored edge case values of the factors. Finally, we validate our methodology in complex SPSys development for the INCARE project (Sec. 6.1) and by robotics practitioners assessment 6.2.
@@
Table 1: Related works comparison with this work
System development method EIV Formal meta-model SCP SE System structure configurability DE E QE MDT Purpose
Digital Mockup [12] - - ES - Single structure - - - Real time (RT) simulation for HIL testing
C2PS [13] - Own meta-model Formal and detailed BS Single adaptive - - N/D Digital Twin in the cloud
[14] Dolev-Yao [15] Formal and detailed - Single structure - - Cyber-security, physical-digital twin synchronisation
DEVSRT [16] - DEVS [17] General - Single structure - - - - Simulation to embedded continuous development
aRD[18] - - General - Multiple setups - - - RT & NonRT parts integration
HMLF [19] - - ES - Single structure - - - - Simulation-driven ML
RSHPN [20] - RSHPN /Petri net ES DI Single structure - - - Control deadlock check, code generation
SPSysML SPSysML /SysML ES / Formal and detailed SPE Multiple setups Structure evaluation
EIV – Embodiment integrity verification, SCP – System creation procedure, SE – Structure evaluation, DE – supports simulated and physical components, DI – Deadlock identification, E – If uncontrolled agents of the physical environment (link humans) can be modelled in the simulation, QE – If quantitative evaluation factors for system design are proposed, MDT – If a multi Digital Twin system model is presented, ES – Example system as the method application procedure, BS – The best selection from the previously designed, SPE – Simulated-Physical design evaluation, N/D – Possible, however, not defined,
@@ 2 Related work
2.1 Literature review constraints
The key related works are described in Table 1, where their crucial features, including goals, are compared with SPSysML. Other works are briefly described in the following sections.
2.2 Simulated-physical system development
In [18], the authors note the robot development process features parallel tasks conducted by heterogeneous R&D teams. The authors suggest system simulation enables rapid prototyping and an iterative development process. They propose the agile Robot Development (aRD) concept facilitating the integration of the hard real-time part and the non-real-time part of the system utilising a Matlab/Simulink toolchain. The sim2real problem is known and mentioned in the recent research results, e.g. in the domain of self-driving cars [21]. However, to our knowledge, none of the existing works specifies a robotic system meta-model with integrity evaluation for mixed simulated-physical setups of CPS.
@@ 2.2 Simulated-physical system development
2.3 System design evaluation
Early-stage product development challenges systems engineers in designing accurate structures and architectures. Evaluating design concepts is crucial as it mightily influence both the quality and cost of the final product.
Tools and models supporting designers in multiattribute utility analysis are the core of this research area. An example Methodology for the Evaluation of Design Alternatives (MEDA) [23] was presented already in 1991. It is based on the subjective designer’s view of the design alternative attributes, e.g. utility. Nowadays, customers play a crucial role in the design process. The authors of [24] tackle the problems that arise in this situation. Priorities are set, and the best design alternative(s) are chosen using the proposed comparison rules. Multi-criteria decision-making is resolved as a constrained multi-attribute optimisation solution [25] or by integrating rough sets in handling vagueness with grey relation analysis [26]. The system’s integrity investigated in our work is another system attribute; thus, our evaluation factors values can provide unbiased, objective input to the design evaluation methods and make the design decision aware of the component reuse level between the system’s simulated and physical embodiments.
@@ 2.4 Meta-models for CPS
Engineers use a specific language to plan, conceptualise, and specify the system. Currently, the state-of-the-art approach is a model-based approach utilising a DSL. There are DSLs supporting verification and testing of, e.g. industry 4.0 plants [27], or agent-based computational systems [28]). Unified Modeling Language (UML) is the most known language; however, System Modeling Language (SysML) [29] is proposed by the Object Management Group (OMG) to support the design, analysis and verification/validation of complex systems comprising software and hardware components. Different Model Driven Engineering approaches exist in the robotics domain [30]. For instance, Embodied Agent-Based cybeR-physical control systems modelling Language (EARL) [31] is SysML specialisation for robotics which allows analysis and specification of the robotic system properties. It is based on the Embodied Agent approach [32]. There are various design analysis tools based on this approach: top-down communication-based system decomposition [33], and Petri net-based design [34] for control deadlocks analysis [20]. It was applied to robotic and multi-modal human-machine interface [35] systems. This article extends the Embodied Agent approach to the simulated and DT-related components domain. We manage this extra complexity with the proposed requirement-based design method and quantitative system design evaluation.
3 Contribution and result validation
@@
None of the models presented in Sec. 2 specifies SPSys in general, especially including a physical part without equivalent Digital Twin, simulated parts without equivalent PT, or hybrid SPSys including Twins and non-Twin components. The key novel features delivered by SPSysML are:
@@
  • Integrity evaluation with quantitative factors based on SysML-based model,
  • Guidance for the simulated-physical integrity increase measured by the quantitative factors,
  • Requirement-based system composition method for straightforward requirement satisfaction tracing,
@@
  • Specification of Simulated-Physical System including multiple physical and simulated parts and multiple Digital Twins,
@@
  • Reflection of the dynamic environment in simulation to enable Digital Twins to observe exogenous actions,
  • A survey-based assessment of Domain-specification Language,
  • Validation of a system design evaluation method in robotics by system implementation in ROS and ROS2.
@@
Based on our experience in SPSys development (TIAGo robot [36, 37, 38], Velma robot [39], IRP6 robot [^1] Real: https://youtu.be/wJpFcy99Gh0, Simulation: https://youtu.be/BjwcbSdouHw [40, 41]) and the literature analysis, we propose the SPSysML allowing software integrity evaluation between the system embodiments, and further by iterative system design, maximise the between embodiments integrity. Complex system structuring is a broad topic, and solutions are utilised using various methods and strategies, e.g., top-down, middle-out, and bottom-up. In [33], we describe a binary communication-focused top-down approach for robotic systems. This article proposes a requirement-based bottom-up system structuring method customised for SPSys.
SPSysML comprises a componet-based Platform Independent Model (PIM), thus, can be applied to any SPSys. As the system is developed, e.g. with Robot Operating Framework (ROS), it becomes a Platform Specific Model (PSM) that can be launched in the specified setups. PSM implementation utilises platform-specific tools and software libraries. Therefore, to validate our approach, we design, implement, test and deploy a specific SPSys. It uses the TIAGo service robot [42] for the INCARE (Integrated Solution for Innovative Elderly Care) project [^2] http://aal-incare.eu/, robot application demonstration (simulated/real) can be found accessing Demo / RAPP Store Demo. We integrate extended voice interface and additional devices to serve the elderly, e.g. in object transportation (Fig. 2(a)) and fall assistance (Fig. 2(b)). The system implementation is component-based and uses frameworks like ROS [43, 44], ROS2 [45], OROCOS [46], that are a standard open-source robot control frameworks. We use two open-source simulators Gazebo [47] (in particular Gazebo_ros_control package [48]) and O3DE [49] for DT implementation. Besides validation by application, we present results of SPSysML assessment among systems engineering practitioners. To make the SPSys development easier for the community, we share the SysML profiles, the SPSysML meta-model and the example INCARE system model [^3] https://github.com/RCPRG-ros-pkg/spsysml. All diagrams describing SPSysML and the INCARE system are also shared [50].
4 Simulated-Physical System